The pH of a 0.14 M solution of an unknown monoprotic acid is 5.85. Calculate the Ka of the acid.
Answer #1

HA <------------> H+ + A-

pH = 5.85

[H+] = 10-pH   = 10-5.85  = 1.41 x 10-6

[A-] = 1.41 x 10-6

[HA] = 0.14 - 1.41 x 10-6  = 0.139998

Ka = [H+] [A-] / [HA]

Ka = [1.41 x 10-6] [1.41 x 10-6] / [0.139998]

Ka = 1.42 x 10-11